A girl is 6 years younger than her oldest brother. The product of their ages is 135. Find their ages.
Answer by corncob(8)

Okay to do this you will need to set up a simple equation and give out variables.
Girls age = x-6
Brother's age= *
the you will take those 2 and add them and place them equil to 135 like so:
x-6+x=135
2x-6+6=135+6
2x=141
2x/2=141/2
x=70.5
The brothers age would be 70.5 then and the girl's would be 64.5